OpenStudy (anonymous):

Indicate which property is illustrated in Step 8. Step 1 2 - 5x - 2 + 6x = (2 - 5x) + (-2 + 6x) Step 2 = (2 - 5x) + (6x - 2) Step 3 = 2 + (-5x + 6x) - 2 Step 4 = 2 + (-5 + 6)x - 2 Step 5 = 2 + x - 2 Step 6 = x + 2 - 2 Step 7 = x + 0 Step 8 = x identity distributive associative commutative @mathmath333

OpenStudy (mathmath333):

what do u think ? we just did this problem

OpenStudy (anonymous):

Np this is a completely different problem, just the same type. Can you give me examples of each of the answer choices?

OpenStudy (mathmath333):

yes wait

OpenStudy (mathmath333):

\(\large\tt \color{blue}{commutative~~property=a+b=b+a}\) \(\large\tt \color{red}{associative~~property=a+(b+c)=(a+b)+c}\) \(\large\tt \color{green}{identity~~property=a+0=a}\) \(\large\tt \color{brown}{distributive~~property=c(a+b)=c\times b+c\times a}\)

OpenStudy (anonymous):

distributive prperty?

OpenStudy (mathmath333):

look at step 8

OpenStudy (anonymous):

identity?

OpenStudy (mathmath333):

yea

OpenStudy (anonymous):

Indicate which property is illustrated in Step 4. Step 1 4 - 11x - 4 + 12x = (4 - 11x) + (-4 + 12x) Step 2 = (4 - 11x) + (12x - 4) Step 3 = 4 + (-11x + 12x) - 4 Step 4 = 4 + (-11 + 12)x - 4 Step 5 = 4 + x - 4 Step 6 = x + 4 - 4 Step 7 = x + 0 Step 8 = x Identity?

OpenStudy (mathmath333):

look at step 4 and 3

OpenStudy (anonymous):

associative?

OpenStudy (mathmath333):

u r taking common x in(-11x + 12x) from step 3 to (-11 + 12)x in step 4

OpenStudy (anonymous):

That's associative right?

OpenStudy (mathmath333):

look at distributive once

OpenStudy (mathmath333):

in distributive pro ,we take c common as x